OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x+2y=8 2x+y=7

OpenStudy (amistre64):

what do you see that we could relate y to x with?

OpenStudy (amistre64):

in other words, solve one of the equations for y or x,

OpenStudy (anonymous):

so like but a number in and solve

OpenStudy (anonymous):

x = 8 - 2y , replace x with this value in the second equation. You get 16 - 4 y +y = 7. so you get y and then x

OpenStudy (amistre64):

not a number so much as just a relationship; we need to be able to express "y" in terms of "x"; or the other way around

OpenStudy (anonymous):

??? can you explain more please i dnt have a teacher for this class and need lots of help

OpenStudy (amistre64):

these equation share a common value for their x and y; we can use that to our advantage by using one to express a relationship between x and y x+2y=8 ; x = 8-2y or 2y = 8 - x we can use this

OpenStudy (amistre64):

take the other equation: 2x+y=7 and use what we know in it: lets use x= 8-2y 2(8-2y) +y = 7

OpenStudy (anonymous):

o so i subtract from both sides

OpenStudy (amistre64):

you algebra it into a relation between x and y; x = this; or y= that

OpenStudy (anonymous):

o ok thx very much for your help

OpenStudy (amistre64):

youre welcome